OpenStudy (anonymous):

John and Alan have a collection of x baseball cards. John has x/4 cards. What fraction of the cards does Alan have? A. x/3 B. 3x/4 C. 3/4 D. 3x

jimthompson5910 (jim_thompson5910):

Let's say Alan had y cards. So because "John and Alan have a collection of x baseball cards" and "John has x/4 cards", we know that x/4+y = x Solve this equation for y to find how many cards Alan has (in terms of x)

OpenStudy (anonymous):

Ummmm 3x?

jimthompson5910 (jim_thompson5910):

close

OpenStudy (anonymous):

3x/4?

jimthompson5910 (jim_thompson5910):

subtract x/4 from both sides to get y = x - x/4

OpenStudy (jackellyn):

x/4 -x = -y 3x/4 = y

jimthompson5910 (jim_thompson5910):

good, if you combine x-x/4, you'll get 3x/4 So y = 3x/4

OpenStudy (anonymous):

thanks guys :)

OpenStudy (anonymous):

i understand